Your little sister is on third the age of your mom. In 12 years, she will be half the age of your mom. How old are your sister and your mom now?
Answer by ikleyn(52776)

.
Let x be the sister age.

Then the mom's age is 3x, according to the condition.


In 12 years the sister will be x+12 years old, while the mom will be 3x+12,

and the condition says that then


2*(x+12) = 3x + 12.


Simplify and solve:


2x + 24 = 3x + 12

24 - 12 = 3x - 2x  ====>  x = 12.


Answer.  The sister is 12 years old;  the mother's age is 3*12 = 36 years.